OpenStudy (idealist10):

Solve 3x(y^2)y'=y^3+x explicitly.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (fateal):

like this???

OpenStudy (anonymous):

\[\begin{align*}3x(y^2)y'&=y^3+x\end{align*}\] Let's try the Bernoulli substitution, \(v=y^3\), so \(v'=3y^2y'\). \[\begin{align*}xv'&=v+x\\\\ v'-\frac{1}{x}v&=1\end{align*}\] The eq. is linear. Find the integrating factor, etc.

OpenStudy (idealist10):

So the integrating factor is 1/x, right?

OpenStudy (anonymous):

Yes

OpenStudy (idealist10):

Got it! Thanks!

OpenStudy (anonymous):

You're welcome!